40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 14:12:03 2022, Yehonathan, Update, BHD, Trying LO phase locking again Screenshot_2022-09-13_17-41-33.png
    Reply  Thu Sep 15 21:12:53 2022, Paco, Update, BHD, LO phase "dc" control lophase_oltf.pdflophase_noise_spectra.pdf
       Reply  Tue Sep 27 10:50:11 2022, Paco, Update, BHD, calibrated LO phase noise lo_phase_asd.pdf
          Reply  Wed Sep 28 16:37:26 2022, Paco, Update, BHD, calibrated LO phase noise; update lo_phase_asd.pdf
             Reply  Wed Sep 28 21:54:08 2022, Paco, Update, BHD, calibrated LO phase noise; update dcpd_phase_asd.pdflo_phase_asd.pdf
                Reply  Fri Sep 30 20:18:55 2022, Paco, Update, BHD, LO phase noise with different actuation points BHDC_asd_actuation_point.png
                   Reply  Mon Oct 3 15:19:05 2022, Paco, Update, BHD, LO phase noise and control after violin mode filters filters_c1sus_lo1_lsc.pngBHDC_asd_act.pngboosted_lo_phase_control.png
Message ID: 17163     Entry time: Wed Sep 28 21:54:08 2022     In reply to: 17161     Reply to this: 17167
Author: Paco 
Type: Update 
Category: BHD 
Subject: calibrated LO phase noise; update 

Repeated the LO phase noise measurement, this time with the LO - ITMY single bounce, and a couple of fixes Koji hinted at including:

  1. The DEMOD angle was the missing piece! The previous error point showed lower noise than the individual DCPDs because the demodulation angle had not been checked. I corrected it so that the error point in LO_PHASE control was exactly equal to the LO-ITMY single bounce fringe. With this, the gain on the servo had to be adjusted from 4.00 to 0.12, still using FM4, FM5, and this time also FM8 (BLP600).
  2. Turned off 60 Hz harmonics comb notches on DCPDs, they are unecessary.
  3. Acquired noise spectra down to 0.1 Hz, with 0.03 Hz bin width to increase resolution and identify resonant SUS noise near 1 Hz.

This time, after alignment the fringe amplitude was 500 counts. Attachment #1 shows the updated plot with the calibrated noise spectra for the individual DCPD signals A and B as well as their rms values. Attachment #2 shows the error point, in loop and the estimated out of loop spectra with their rms as well. The peak at ~ 240 Hz is quite noticeable in the error point time series, and dominates the high frequency rms noise. The estimated rms out of loop noise is ~ 9.2 rad, down to 100 mHz.

Attachment 1: dcpd_phase_asd.pdf  243 kB  Uploaded Wed Sep 28 23:28:09 2022  | Hide | Hide all
dcpd_phase_asd.pdf
Attachment 2: lo_phase_asd.pdf  243 kB  Uploaded Wed Sep 28 23:28:46 2022  | Hide | Hide all
lo_phase_asd.pdf
ELOG V3.1.3-